1/1

Arabella Katherine Hankey,

Tell me the old, old story

Tell me the old, old story of unseen things above, Of Jesus and His glory, of Jesus and His love. Tell me the story simply, as to a little child, For I am weak and weary, and helpless and defiled. Tell me the old, old story, tell me the old, old story, Tell me the old, old story, of Jesus and His love.

Tell me the story slowly, that I may take it in, That wonderful redemption, God’s remedy for sin. Tell me the story often, for I forget so soon; The early dew of morning has passed away at noon. Tell me the old, old story, tell me the old, old story, Tell me the old, old story, of Jesus and His love.

Tell me the story softly, with earnest tones and grave; Remember I’m the sinner whom Jesus came to save. Tell me the story always, if you would really be, In any time of trouble, a comforter to me. Tell me the old, old story, tell me the old, old story, Tell me the old, old story, of Jesus and His love.

Tell me the same old story when you have cause to fear That this world’s empty glory is costing me too dear. Yes, and when that world’s glory is dawning on my soul, Tell me the old, old story: “Christ Jesus makes thee whole.” Tell me the old, old story, tell me the old, old story, Tell me the old, old story, of Jesus and His love.

Verkinto de tiu ĉi Angla poemo estas Arabella Katherine Hankey (∗1834 – †1911).

Arg-1026-2080 (2014-05-28 09:48:31)

Pri la verkintino vidu la retejon http: // cyberhymnal. org/ bio/ h/ a/ n/ hankey_ak. htm . La melodio estas verkita de William H. DOANE, en 1870. Legu ankaŭ: http: // www. reocities. com/ cigneto/ thctxt/ d/ dirulahis1. html .

http://www.poezio.net/version?poem-id=1026&version-id=,2080